Changes between Version 16 and Version 17 of Status


Ignore:
Timestamp:
Jul 15, 2008 3:45:26 PM (11 years ago)
Author:
gaew
Comment:

--

Legend:

Unmodified
Added
Removed
Modified
  • Status

    v16 v17  
    44
    55Please contact Gary Llewellyn (Science / operations coordinator) if you have any questions or concerns: gaew@nerc.ac.uk (01865 374 391)
     6
     7== 15th July ==
     8''Yesterday (14.07.08) a failure in the AISA Eagle control system cause us to abort an attempt to collect data from the North bank of the Thames; this problem has been corrected. Unfortunately, despite the present good weather in the South and East of the UK conditions continue to be unfavourable in the Midlands, Cumbria and Scotland. We hope to be able to collect data from Wytham Woods (GB08-08) and the Thames (HY05-04) this week, however, the passive remote sensing instruments are due to be removed from the aircraft on Friday (18.07.08) in preparation for calibration/validation with FSF in following week. The research teams for one of the Peak District (GB08-02) projects and the project based in North Wales (GB08-13) have deferred the next collection for these projects to August when data collection (with the LiDAR fitted on DCALM) should resume on the 12th of August (12.08.08).''
    69
    710== 9th July ==